Greg is 4 times older than stab. 8 years from now Greg will be twice as old as stab. How old will stab be then?
Answer by Maths68(1474)

Let
Present age of Greg = g
Present age of Stab = s

Greg is 4 times older than stab.
g=4s............(1)
8 years from now
Greg will be = g+8
Stab will be = s+8
Then
Greg will be twice as old as stab.
g+8=2(s+8)
g+8=2s+16
g=2s+16-8
g=2s+8...........(2)
Put the value of g from (1) to (2)
g=2s+8
4s=2s+8
4s-2s=8
2s=8
s=8/2
s=4
Put the value of s in (1)
g=4s
g=4*4
g=16

Present age of Greg = g = 16 years old
Present age of Stab = s = 4 years old
8 Years from Now
Greg will be = 16+8 = 24 years old
Stab will be = 4+8 = 12 years old